# EulerFormer: Sequential User Behavior Modeling with Complex Vector Attention

### v1

@@ Abstract.
To address this issue, this paper proposes a novel transformer variant with complex vector attention, named EulerFormer, which provides a unified theoretical framework to formulate both semantic difference and positional difference. The EulerFormer involves two key technical improvements. First, it employs a new transformation function for efficiently transforming the sequence tokens into polar-form complex vectors using Euler’s formula, enabling the unified modeling of both semantic and positional information in a complex rotation form. Secondly, it develops a differential rotation mechanism, where the semantic rotation angles can be controlled by an adaptation function, enabling the adaptive integration of the semantic and positional information according to the semantic contexts. Furthermore, a phase contrastive learning task is proposed to improve the anisotropy of contextual representations in EulerFormer. Our theoretical framework possesses a high degree of completeness and generality (e.g., RoPE can be instantiated as a special case). It is more robust to semantic variations and possesses more superior theoretical properties (e.g., long-term decay) in principle. Extensive experiments conducted on four public datasets demonstrate the effectiveness and efficiency of our approach. Our code is available at https://github.com/RUCAIBox/EulerFormer.
@@ 4. EXPERIMENTS
To verify the effectiveness of the proposed EulerFormer, we conduct extensive experiments and report detailed analysis results.
@@ 4.1.1. Datasets and Evaluation Metrics
To evaluate the performance of EulerFormer, we conduct experiments on four publicly available datasets: (1) MovieLens (Harper and Konstan, 2015) contains users’ ratings on movies, which is a popular dataset for sequential recommendations. We adopt two widely used versions: ML-1M and ML-20M. (2) Yelp is a business dataset containing user reviews for various restaurants. We adopt its latest version, which was released in 2022 (Yelp2022). (3) Amazon_Books (McAuley et al., 2015) is part of the Amazon review datasets, covering rich user interactions with an extensive range of books. Following existing work (Fan et al., 2021; Hou et al., 2022b, 2023), we group user interactions in chronological order across all datasets and split the data using the leave-one-out strategy. Moreover, we use three common evaluation metrics, including Recall, MRR, and NDCG, for the evaluation of top-$k$ ($k=10$) recommendations. We rank the ground-truth item of each sequence among all other items for evaluation on the test set and finally report the average score of all test users.
@@
To show the effectiveness of our model, we integrate EulerFormer into existing state-of-the-art sequential recommenders, including: (1) SASRec (Kang and McAuley, 2018) adopts the transformer for obtaining the sequence representation. (2) BERT4Rec (Sun et al., 2019) adapts the BERT (Devlin et al., 2018) model by using the cloze task to enhance sequence representations. (3) CORE (Hou et al., 2022a) proposes to unify the representation space for generating better representations. (4) SASRecF (Kang and McAuley, 2018) improves SASRec (Kang and McAuley, 2018) by introducing item features as additional contextual information. (5) FDSA (Zhang et al., 2019) equips multiple transformers to learn item-wise and feature-wise representations. Further, we also compare our model with the state-of-the-art positional encoding methods in NLP, including: (1) Sinusoidal (Vaswani et al., 2017) uses sinusoidal functions to generate positional embeddings. (2) XLNet (Dai et al., 2019) explicitly employs relative position embeddings to learn relative position information. (3) RoPE (Su et al., 2024) sets specific rotatory matrices based on the position of each key or query, enabling the computation of attention scores using relative positional information.
@@
We implement EulerFormer using a popular recommendation library RecBole (Zhao et al., 2021). We use the Adam (Kingma and Ba, 2014) optimizer and carefully search the hyper-parameters of the compared methods. The dimension of the latent vectors is fixed to 64. Following the original papers, we set the hyper-parameters of Transformer for all sequential models, including the number of layers as $L=2$, attention head as $h=2$, inner size of feed-forward layers as 256, and dimension size as $d=64$. For EulerFormer, we search the temperature parameters $\tau$ (See in Eq. (14)) among $\{0.5,1,5,10\}$ and loss weight $\epsilon$ (See in Eq. (18)) among $\{1e-4,1e-5,1e-6\}$.
@@ 4.2. Overall Performance
To demonstrate the effectiveness of EulerFormer in improving recommendation performance, we integrate it into representative Transformer-based recommender models. We report the main experimental results in Table 3 and make the following observations:
(1) For all methods, equipped with EulerFormer, the model performs better than the original ones in all cases. Notably, integrating EulerFormer into SASRec (Kang and McAuley, 2018) (EulerFormer+SASRec) exhibits improvements of 10.46% and 8.30% with respect to the NDCG metric on ML-1M and Yelp2022 dataset respectively. It demonstrates the effectiveness of our approach for enhancing the positional encoding capacity of transformer architecture.
@@
(3) Our method shows a smaller improvement for SASRecF (Kang and McAuley, 2018) and FDSA (Zhang et al., 2019) than for SASRec (Kang and McAuley, 2018). It may be caused by the fact that item features will enhance the isotropy of the representations, leading to the model being less sensitive to positional information.
Table 3. Performance (%) of backbone models (w): use EulerFormer and (w/o): use original Transformer. All improvements are statistically significant (i.e., two-sided $t$-test with p ¡ 0.01) over original models.
@@
To demonstrate the superiority of EulerFormer over the state-of-the-art positional encoding models (See in Section 4.1.2), we take the SASRec as the base model and report the experimental results in Table 4. Accordingly, we have the following observations:
(1) XLNet (Dai et al., 2019) outperforms Sinusoidal (Vaswani et al., 2017) on the Yelp2022, Amazon_Books and ML-20M datasets, which shows the superiority of relative positional encoding in handling sparse datasets.
@@
As for the efficiency, we can observe that the base model (learned position embedding (Vaswani et al., 2017)) is the most efficient since its structure is simple and does not require additional calculations. Compared to Sinusoidal (Vaswani et al., 2017), RoPE (Su et al., 2024) requires the reversal of embedding dimensions, and thus it has higher latency. The latency of XLNet (Dai et al., 2019) is much larger due to the complicated positional encoding strategy. As a comparison, the latency of EulerFormer is much less than RoPE (Su et al., 2024) and XLNet (Dai et al., 2019), and it is comparable to the efficient algorithm Sinusoidal (Vaswani et al., 2017). With the highest accuracy and linear complexity, EulerFormer has a great potential to be applied into large-scale recommender systems.
@@ 4.3. Further Analysis of EulerFormer
In this section, we further perform a series of detailed analysis on the proposed EulerFormer to confirm its effectiveness. Due to the limited space, we only report the results on ML-1M and Yelp2022 datasets, and the observations are similar on other datasets.
@@
We analyze how our proposed components influence the final encoding performance. We prepare four variants of the proposed EulerFormer model for comparisons, including: (1) w/o adaptation function that removes the weights $\bm{\delta}$ and bias $\bm{b}$ in Eq. (13), i.e., directly uses Eq. (10) for positional encoding; (2) w learnable positional encoding that removes the adaptive weight $\bm{\delta}$ and $\bm{b}$ and sets $\bm{g}$ in Eq. (10) as a learnable parameter; (3) w/o differential rotation that removes the relative positional encoding module (i.e., $\Delta\bm{P}$ in Eq. (10)); (4) w/o rotary embedding that removes the rotary embedding in Eq. (5); (5) w/o phase-contrastive learning that removes the phase contrastive learning in the training process. The results are shown in Table 5. We can observe that all the proposed methods are useful to improve the performance. Specifically, the variant (1) and (2) show a great decrease, which indicates that our proposed adaptive rotation mechanism is more suitable to model the positional information in the self-attention formulation.
@@ 4.3.3. Analysis of Adaptive Rotations
To further understand how EulerFormer helps the models deal with sequential data, we conduct a case study to visualize the attention scores. In Figure 4, we show a user’s attention score map from the ML-1M dataset, where the lighter the color is, the higher the attention score is. Comparing the heatmaps generated by SASRec and SASRec+EulerFormer, we can observe that the attention of SASRec is relatively dispersed, hindering its ability to effectively capture useful information from the sequence. Equipped with EulerFormer, SASRec tends to assign larger attention scores on recent items, which indicates the improved capacity of capturing the relative position information of the sequential patterns. This is consistent with prior work (Su et al., 2024). It indicates the properties of EulerFormer in reducing the inter-token dependency as relative distances increase, which enables the model to focus on tokens that are closer in distance, thereby making it more capable of handling long-sequence data.
@@
As introduced in Section 3.3, we introduce a phase-contrastive task to enhance the anisotropy between different tokens. To verify this, we visualize the phase distribution of sequence tokens at different positions on the ML-1M dataset. As shown in Figure 5, for ease of illustration, we draw the phases located at different positions on different circular orbits, and the phases are displayed on those circles. We can observe that, without the phase-contrastive learning task, the phase distributions at different positions are very close. After undergoing phase contrastive learning, the differences in phase distributions at different positions become significant. These results demonstrate that our method can effectively enhance the anisotropy of item representations at different positions. This enhancement enables the model to effectively distinguish items at various positions, thereby improving the effectiveness of positional encoding.
@@ 4.3.6. Impact of the Coefficient $\epsilon$
In the phase contrastive learning loss defined in Eq. (18), the coefficient $\epsilon$ can balance the two losses for sequential recommendations. To analyze the influence of $\epsilon$, we vary $\epsilon$ from 5e-6 to 3e-5 and report the results in Figure 6(a). It demonstrates that an appropriate $\epsilon$ can effectively improve the recommendation performance of the EulerFormer. Specifically, setting the hyper-parameter $\epsilon$ to approximately 1e-5 results in enhanced performance on both ML-1M and Yelp2022 datasets.
@@ 4.3.7. Impact of the Temperature $\tau$
To analyze the impact of temperature (See in Eq. (14)) on EulerFormer, we vary $\tau$ in the range of 0.6 to 1.6 and show the results in Figure 6(b). We can observe that too large a value of $\tau$ will cause poor performance, which is consistent with the experimental results reported in prior work (You et al., 2020; Lin et al., 2022). In addition, the optimal temperature for the Yelp2022 dataset is lower, indicating that the temperature setting for EulerFormer should be reduced for more sparse datasets. Generally, a temperature in the range of 0.8 to 1.2 can lead to good recommendation performance.

### v2

@@ Abstract.
To address this issue, this paper proposes a novel transformer variant with complex vector attention, named EulerFormer, which provides a unified theoretical framework to formulate both semantic difference and positional difference. The EulerFormer involves two key technical improvements. First, it employs a new transformation function for efficiently transforming the sequence tokens into polar-form complex vectors using Euler’s formula, enabling the unified modeling of both semantic and positional information in a complex rotation form. Secondly, it develops a differential rotation mechanism, where the semantic rotation angles can be controlled by an adaptation function, enabling the adaptive integration of the semantic and positional information according to the semantic contexts. Furthermore, a phase contrastive learning task is proposed to improve the isotropy of contextual representations in EulerFormer. Our theoretical framework possesses a high degree of completeness and generality (e.g., RoPE can be instantiated as a special case). It is more robust to semantic variations and possesses more superior theoretical properties (e.g., long-term decay) in principle. Extensive experiments conducted on four public datasets demonstrate the effectiveness and efficiency of our approach. Our code is available at https://github.com/RUCAIBox/EulerFormer.
@@
GSAI is the abbreviation of Gaoling School for Artificial Intelligence
@@ 4.1.1. Datasets and Evaluation Metrics
We conduct experiments on four publicly available datasets: (1) MovieLens contains users’ ratings on movies. We adopt two widely used versions: ML-1M and ML-20M. (2) Yelp contains user reviews for various restaurants. We adopt its latest version (Yelp2022). (3) Amazon_Books covers rich user behaviors with an extensive range of books. We split the data using the leave-one-out strategy and use three common evaluation metrics, i.e., Recall, MRR, and NDCG, for the evaluation of top-$10$ recommendations. We rank the ground-truth item among other items on the test set and report the average score of all users.
@@
To show the effectiveness of our model, we integrate EulerFormer into several representative sequential models, including: (1) SASRec (Kang and McAuley, 2018) adopts the transformer for obtaining the representation. (2) BERT4Rec (Sun et al., 2019) uses the cloze task to enhance sequence representations. (3) CORE (Hou et al., 2022a) unifies the representation space for generating better representations. (4) SASRecF (Kang and McAuley, 2018) improves SASRec (Kang and McAuley, 2018) by using item features as additional information. (5) FDSA (Zhang et al., 2019) uses multiple transformers to learn item-wise and feature-wise representations. Further, we also compare our model with the state-of-the-art positional encoding methods in NLP, including: (1) Sinusoidal (Vaswani et al., 2017) uses sinusoidal functions to generate positional embeddings. (2) XLNet (Dai et al., 2019) employs relative embeddings to learn relative positional relations. (3) T5 bias (Raffel et al., 2020) uses attention bias to model relative positions. (4) ALiBi (Press et al., 2021) uses specific positional bias to model pair-wise positions. (5) RoPE (Su et al., 2024) uses specific rotatory matrices to model the positional information.
@@
We implement EulerFormer using a popular recommendation library RecBole (Zhao et al., 2021, 2022; Xu et al., 2023). We use the Adam (Kingma and Ba, 2014) optimizer and carefully search the hyper-parameters of the compared methods. The dimension of the embeddings vectors is fixed to 64. Following the original papers, we set the hyper-parameters of Transformer for all sequential models, including the number of layers as $L=2$, attention head as $h=2$, inner size of feed-forward layers as 256, and dimension size as $d=64$. For EulerFormer, we search the temperature parameters $\tau$ among $\{0.5,1,5,10\}$ and loss weight $\epsilon$ among $\{1e-4,1e-5,1e-6\}$.
@@ 4.2. Overall Performance
To show the effectiveness of EulerFormer in improving recommendation performance, we integrate it into representative sequential recommenders. As shown in Table 3, we have the following observations:
(1) For all methods, equipped with EulerFormer, the model performs better than the original ones in all cases. It shows the effectiveness of our approach for enhancing the positional encoding capacity of transformers.
@@
(3) Our method shows a smaller improvement for SASRecF (Kang and McAuley, 2018) and FDSA (Zhang et al., 2019) than for SASRec (Kang and McAuley, 2018). It may be caused by the fact that item features will enhance the anisotropy of the representations, leading to the model being less sensitive to the position.
Table 3. Performance (%) of backbone models (w): use EulerFormer and (w/o): use original Transformer. The best result is bold, while the second-best result is underlined. “Improv.” indicates the relative improvement ratios of the EulerFormer over the original Transformer.
@@
To show the superiority of EulerFormer over the SOTA positional encoding models, we take the SASRec as the base model and report the results in Table 4. Accordingly, we have the following observations:
(1) Relative positional encoding models (i.e., XLNet (Dai et al., 2019), ALiBi (Press et al., 2021) and T5 bias (Raffel et al., 2020)) outperform Sinusoidal (Vaswani et al., 2017) on the Yelp2022, Amazon_Books and ML-20M datasets, which shows the superiority of learning relative positional relations in handling sparse datasets.
@@
For efficiency, ALiBi (Press et al., 2021), T5 bias (Raffel et al., 2020) and the base model (Vaswani et al., 2017) are more efficient since their structures are simple and do not require additional calculations. Compared to Sinusoidal (Vaswani et al., 2017), RoPE (Su et al., 2024) requires the reversal of embedding dimensions, and thus it has higher latency. The latency of XLNet (Dai et al., 2019) is much larger due to the complicated positional encoding strategy. As a comparison, the latency of EulerFormer is much less than XLNet (Dai et al., 2019), and it is comparable to the efficient algorithm Sinusoidal (Vaswani et al., 2017). With the highest accuracy and linear complexity, EulerFormer has a great potential to be applied into large-scale recommender systems.
@@ 4.3. Further Analysis of EulerFormer
We analyze how our proposed components influence the final encoding performance. We prepare four variants for comparisons including: (1) w/o adaptation function that removes the weights $\bm{\delta}$ and bias $\bm{b}$ in Eq. (13); (2) w learnable positional encoding that removes the adaptive weight $\bm{\delta}$ and $\bm{b}$ and sets $\bm{g}$ in Eq. (10) as a learnable parameter; (3) w/o differential rotation that removes the relative positional encoding module (i.e., $\Delta\bm{P}$ in Eq. (10)); (4) w/o rotary embedding that removes the rotary embedding in Eq. (5); (5) w/o phase-contrastive learning that removes the phase contrastive learning in the training process. The results are shown in Table 5. All the proposed methods are useful to improve the performance. Specifically, the variant (1)(2) show a great decrease, which indicates that our proposed adaptive rotation mechanism is more suitable to model the positional information in the self-attention formulation.
@@ 4.3.3. Analysis of Adaptive Rotations
To further understand how EulerFormer helps the models deal with sequential data, we conduct a case study to visualize the attention scores. In Figure 4, we show a user’s attention map (the 1st head of the 1st layer) from the ML-1M dataset, where the lighter the color is, the higher the attention score is. Comparing the heatmaps generated by SASRec and SASRec+EulerFormer, we can observe that the attention of SASRec is relatively dispersed, hindering its ability to capture useful information from the sequence. Equipped with EulerFormer, SASRec tends to assign larger attention scores on recent items, which indicates the improved capacity of capturing the relative position information of the sequential patterns. This is consistent with the prior work (Su et al., 2024). It indicates the properties of EulerFormer in reducing the inter-token dependency as relative distances increase, which enables the model to focus on tokens that are closer in distance, thereby making it more capable of handling long-sequence data.
@@
As introduced in Section 3.3, we introduce a phase-contrastive task to enhance the isotropy between different tokens. To verify this, we use the SASRec as the base model and visualize the phase distribution of sequence tokens at different positions after training on the ML-1M dataset. As shown in Figure 5, for ease of illustration, we draw the phases (See Eq. (3)) located at different positions on different circular orbits, and the phases are displayed on those circles. We can observe that, when removing the phase-contrastive learning task, the phase distributions at different positions are very similar. After undergoing phase-contrastive learning, the differences in phase distributions at different positions become significant. These results demonstrate that our approach can effectively enhance the isotropy of item representations at different positions. This enables the model to effectively distinguish items at various positions, thereby improving the effectiveness of positional encoding and sequence modeling.
@@ 4.3.6. Impact of the Coefficient $\epsilon$
In the phase contrastive learning loss defined in Eq. (18), the coefficient $\epsilon$ can balance the two losses for sequential recommendations. To analyze the influence of $\epsilon$, we vary $\epsilon$ from 5e-6 to 3e-5 and report the results (w.r.t. SASRec) in Figure 6(a). It demonstrates that an appropriate $\epsilon$ can effectively improve the recommendation performance of the EulerFormer. Specifically, setting the hyper-parameter $\epsilon$ to approximately 1e-5 results in enhanced performance on both ML-1M and Yelp2022 datasets.
@@ 4.3.7. Impact of the Temperature $\tau$
To analyze the impact of temperature (See in Eq. (14)), we vary $\tau$ in the range of 0.6 to 1.6 and show the results (w.r.t. SASRec) in Figure 6(b). We can observe that too large a value of $\tau$ will cause poor performance, which is consistent with the experimental results reported in prior work (You et al., 2020; Lin et al., 2022). In addition, the optimal temperature for the Yelp2022 dataset is lower, indicating that the temperature setting for EulerFormer should be reduced for more sparse datasets. Generally, a temperature in the range of 0.8 to 1.2 can lead to good recommendation performance.